OpenStudy (anonymous):

A spring will stretch 12 cm when a 250-g mass is hung on it. If two springs are hung side by side and both support the 250-g mass, how much will each spring stretch? Explain.

OpenStudy (doc.brown):

If you share the force of the mass being pulled by gravity evenly, you cut the force in half. \[F=ma\]\[\frac{F}{2}=\frac{ma}{2}\] The force by the spring is the same. \[F=kx\]\[\frac{F}{2}=\frac{kx}{2}\] or \[\frac{ma}{2}=\frac{kx}{2}\] The spring constant k will remain... um... constant. So the distance x will be halved.

OpenStudy (anonymous):

@doc.brown can you please explain it more? i understand F=MA but i dont understand what numbers to put where in order to get the answer. i know the mass is 250-g because it's stated in the problem.

OpenStudy (anonymous):

correct m = 250g, and inorder to put it into the equation it needs to be in kg, so 250g becomes 0.250kg ^_^ as for the other variables: a = acceleration x = distance the spring is stretched k = 'spring constant' does this help? @colbiespann

OpenStudy (anonymous):

okay so we will be using F=ma..... sooo... M= 0.250kg x= 12cm so far that's all i have, how do i figure out K and A? @DemolisionWolf

OpenStudy (anonymous):

good, yes you are right I'll help you out with a, a actually stands for acceleration which in our case will be gravity. sooo, a = 9.81 ^_^ now i'm going to approach it a bit different then the guy above...

OpenStudy (anonymous):

so here is what the first thing looks like